Sol Summary – January 06th


Crew 219 Sol Summary Report 06-01-2020

Sol: 1

Summary Title: "I have 5 high-def pictures of a toilet. Awesome!"

Author's name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Sol 1 began with inclement weather and EVA delays at the Commander's discretion. Instead, Crew 219 spent the morning polishing our mission plan, conducting maintenance tasks around campus, reorganizing and consolidating sundries, and preparing materials for scientific work. After a lunchtime reevaluation of the weather, both EVAs were cleared to go, and both were conducted to standard. The late afternoon consisted of sample processing in the science dome, herb harvesting in the GreenHab, report writing in the hab, and the successful execution of a toilet procedure provided by the Director.

Look Ahead Plan: Tomorrow holds plans for two proposed sample collection EVA's, maintenance on a couple EVA suits, and continued sample processing in the science dome.

Anomalies in work: None

Weather: Cold and foggy with intermittent light flurries in the early morning, to cold and overcast in the afternoon, with temperatures still holding just above freezing.

Crew Physical Status: 1 minor infury (Non-EVA) annotated in today's HSO report

EVA: EVA #1 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)
EVA #2 - Terrain Familiarization and Sample Collection at Marble Ritual (0518700/4250700)

Reports to be filed:

Sol Summary
EVA Reports 1 & 2
EVA Requests 3 & 4
Operations Report
HSO Report
GreenHab Report
Research Report
Journalist's Report
Mission Plan
Updated HSO Pre-Mission Checklist
Photos (Including Photo of the Day)

Support Requested:

Sundries: Paper towels
Medical: C-Spine needs to be replaced (broken/misshapen)
12x Medium Gauze Pads
Anitbiotic Ointment

Copyright © The Mars Society. All rights reserved. | Main Site